A ball is projected from ground level at an initial angle of 33.1 degrees above the horizontal. After

0.74 seconds, the ball hits a wall that is 28.4 meters away. At what height above the ground does

the ball hit the wall?



1
Expert's answer
2021-05-24T18:03:09-0400
"28.4=v\cos{33.1}(0.74)\\v=45.8\frac{m}{s}\\h=45.8(0.74)\sin{33.1}-0.5(9.8)(0.74)^2\\h=15.8\ m"